OpenStudy (anonymous):

solve for x, a) 6*2^x= 192 b) 4*(1/3)^x=324

OpenStudy (anonymous):

a) 6*2^x= 192 divide both sides by 6 2^x = 32 2^x = 2^5 Since LHS = RHS and the bases are same (i.e.) 2, hence for equality the powers also have to be same therefore x = 5

OpenStudy (angela210793):